OpenStudy (anonymous):

Some birds and squirrels were on a branch of a tree. There were 15 animals on the branch with a total of 50 animal feet. How many birds and how many squirrels were on the branch? Part 1: Set up the problem. (2 points) Part 2: Explain why you set it up that way. (1 point) Part 3: Provide the solution. (1 point)

OpenStudy (anonymous):

First you have to set up how many feet there are: x= birds feet ; y= squirrel feet. 2x+4y=50

OpenStudy (anonymous):

Then you have to give values to X and Y in order to solve the ecuation.

OpenStudy (anonymous):

x + y = 15 2x +4y = 50

OpenStudy (anonymous):

Solve the system of equations...substitution? y =15-x 2x + 4(15-x) = 50 2x + 60 -4x = 50 -2x + 60 = -10 -2x = -10 x = 5 5 birds, 10 squirrels

OpenStudy (anonymous):

x + y = 15 squirrels and birds = 15 animals 2x + 4y = 50 birds have two legs, squirrels have four legs y = 15 - x solved for y so I could substitute into the second equation

OpenStudy (anonymous):

So why did you set it up the way it is?

OpenStudy (anonymous):

I needed more information than just the number of feet per animal. I could have figured out the solution, but it was faster knowing there were 15 animals.

OpenStudy (anonymous):

The problem gave you two conditions. 1. there were 15 birds and squirrels 2. the feet added together totaled 50

OpenStudy (anonymous):

Thank you for your help!